ab2 and ab4 when dissolved in 20gms of benzene 1 gm of ab2 delta Tf by 2.4k where on 1gm ab4

Grade:12

1 Answers

Arun
3 years ago
As We know that:
 
MB = (Kf x wB x 1000) / (wA X ΔTf)
 
Now ΔTf = 2.3 , wB = 1.0 , wA = 20, KF = 5.1 (given)
 
PUTTING THE VALUES IN THE EQUATION
 
MB = (5.1 x 1 x 1000) / (20 x 2.3) = 110.87 g/mol
 
Therefore MAB2 = 110.9
 
For AB4 compound
 
ΔTf = 1.3 , wb = 1 ,wa = 20
 
MB = (5.1 X 1 X 1000) / (20 X 1.3) = 196 g/mol
 
Therefore MAB = 196
 
Let x be the atomic mass of A & y be the atomic mass of B,
 
THEN MAB2 = x + 2y = 110.9 ----------------------------------(1)
 
And MAB = x + 4y = 196        ----------------------------------(2)
 
Subtracting 2 from 1 ,we get
 
2y  = 196-110.9
 
y   = 85.1 / 2 
 
y = 42.6
 
Putting the value of y in 1 we get
 
x = 110.9 - 2 x 42.6
 
x = 25.59
 
Therefore atomic mass of A = 25.59 u Atomic mass of B = 42.6 u.